OpenStudy (tia_sweetheart17):

find the original price of a pair of shoes if the sale price is $11 after a 75% discount

OpenStudy (anonymous):

11$ = 75% of x 75% can be represented as 75/100 = 0.75 algebraically: 11 = 0.75x solve for x

OpenStudy (tia_sweetheart17):

i have no idea

OpenStudy (anonymous):

divide both sides by 0.75 \[11 = 0.75x\]\[\frac{ 11 }{ 0.75 } = x\]you'll need a calculator for this

OpenStudy (anonymous):

not NEED absolutely but it's much faster :P

OpenStudy (anonymous):

11/0.25 = $44